Aries:

The circumference of a circle measures 11.27xpie ft. What is the measure of the diameter of this circle?

Luigi0210:

The circumference of a circle can be defined as \[\large C= 2 \pi r \] Since they gave you the circumference of 11.27 pi, just plug it in and solve: \[\large 11.27 \pi = 2\pi r\] Divide by 2 pi \[\frac{11.27\pi}{2\pi}= r \] The pi's can cancel, leaving \[\large \frac{11.27}{2} = r \] Since radius is just HALF the diameter, you could multiple by 2 \[\large 11.27 = 2r=D\]

Aries:

okay!